# Differentially Private Matchings

### v2

@@ 3 Lower Bound for Explicit Matchings
Fix some graph $G\in\mathcal{M}$, and let $\{u,v\},\{u^{\prime},v^{\prime}\}\in E(G)$. Let $G^{\prime}$ be the graph obtained by removing $\{u,v\}$ and $\{u^{\prime},v^{\prime}\}$, and adding $\{u,v^{\prime}\}$ and $\{u^{\prime},v\}$. Note that $G$ and $G^{\prime}$ are distance $2$ away from each other.
For all $v^{\prime\prime}\neq u,v^{\prime}$, it must be the case that $p(\mathcal{A},G^{\prime},\{u,v^{\prime\prime}\})\leq b/(n-2)$, since otherwise the symmetry property of Lemma 3.1 and linearity of expectations implies that the expected degree of $u$ in the output of $\mathcal{A}(G)$ is at least $\sum_{v^{\prime\prime}\neq u,v^{\prime}}p(\mathcal{A},G,\{u,v^{\prime\prime}\})>(n-2)\frac{b}{n-2}=b$, contradicting our assumption that $\mathcal{A}$ always outputs a $b$-matching. Thus $p(\mathcal{A},G^{\prime},\{u,v\})\leq b/(n-2)$. Since $G$ and $G^{\prime}$ are at distance $2$ from each other, this implies that $p(\mathcal{A},G,\{u,v\})\leq e^{2\varepsilon}\frac{b}{n-2}+2\delta$.
Since this argument applies to all $\{u,v\}\in E(G)$, we have that
$\text{E}[\nu_{\mathcal{A}}(G)]$ =$=\sum_{\{u,v\}\in E(G)}p(\mathcal{A},G,\{u,v\})$
$\leq\sum_{\{u,v\}\in E(G)}\left(e^{2\varepsilon}\frac{b}{n-2}+2\delta\right)$
$\leq e^{2\varepsilon}\frac{bn}{2(n-2)}+2\delta\frac{n}{2}.$
$\leq e^{2\varepsilon}b+\delta n.$
Since $\nu_{\mathcal{A}}=\min_{G\in\mathcal{M}}\text{E}[\nu_{\mathcal{A}}(G)]$, this implies that $\nu_{\mathcal{A}}\leq e^{2\varepsilon}b+\delta n$, as claimed. ∎
@@
The combination of Lemma 3.1 and Lemma 3.2 imply that any $(\varepsilon,\delta)$-DP algorithm $\mathcal{A}$ which always outputs a $b$-matching has $\nu_{\mathcal{A}}\leq e^{2\varepsilon}b+\delta n$. We also know that $\nu(G)=n/2$ for all $G\in\mathcal{M}$. Thus by the definition of an $(\alpha,\beta)$-approximation, it must be the case that $\alpha(e^{2\varepsilon}b+\delta n)+\beta\geq n/2$. ∎
4 $\varepsilon$-Local Edge Differentially Private Implicit Matchings
@@
In this section, we state our main algorithm for maximal matching (Algorithm 2) and prove its guarantees in Theorem 1.3, which is restated below for readibility. We modify Algorithm 2 in the later sections for various different models. See 1.3
@@
Our LEDP algorithm is based on a simple procedure for maximal non-private $b$-matching, described as follows. We take an arbitrary ordering on the vertices, and process them one by one. When considering the $i^{th}$ vertex $v_{i}$, let $b^{\prime}$ be the number of additional vertices $v_{i}$ can match with. We wish to find some subset of vertices of size $b^{\prime}$ from later in the ordering with which to match $v_{i}$. Some of these vertices may have already matched with $b$ vertices from previous iterations. We choose an arbitrary subset of $b^{\prime}$ vertices which have not already reached their limit from later in the ordering. We say that vertices which have reached their limit satisfy the matching condition. If there are fewer than $b^{\prime}$ vertices satisfying the matching condition, we match all of them with $v_{i}$. Then we move on to the next vertex in the order. It is clear that this procedure yields a maximal $b$-matching. In the following algorithms, we analyze a suitable privatization of this algorithm and show that given sufficiently large $b$, we are guaranteed to produce an implicit maximal ($1$-)matching. [^7] We remark that all of our algorithms can be straightforwardly adjusted to guarantee maximal $b$-matchings.
High-Level Overview
@@ 4.1 Detailed Algorithm Description
We now describe our algorithm, Algorithm 2, in detail. The algorithm is provided with a private graph $G=(V,E)$, a privacy parameter $\varepsilon>0$, a matching parameter $b\geq\frac{576c\ln(n)}{\eta\cdot\varepsilon}+1$ and a constant $c\geq 1$ which is used in the high probability guarantee. We first set the variables $\eta$ and $\varepsilon^{\prime}$ that we use in our algorithm (Algorithms 2 and 2). Then, we flip our public set of coins. We flip $O(\log(n))$ coins for each pair of vertices $u\neq v\in V\times V$ (Algorithm 2). Each of the coins flips for pair $\{u,v\}$ is flipped with probability $1/(1+\eta)^{r}$ for each $r\in\{0,\dots,\lceil\log_{1+\eta}(n)\rceil\}$. The result of the coin flip is released and stored in $coin(u,v,r)$ (Algorithm 2).
@@
Input: Graph $G=(V,E)$, privacy parameter $\varepsilon>0$, matching parameter $b\geq\frac{576c\ln(n)}{\varepsilon}+1$, and constant $c\geq 3$
@@ 9       $A(u)\leftarrow\infty$
12       for each node $u\in V$ such that $A(u)>i$  do
             // $u$ has not satisfied its matching condition
@@ 15                   $u$ releases $A(u)\leftarrow i$
18       if $A(v)>i$  then
             // $v$ has not satisfied its matching condition
@@ 28                   $M(v)\leftarrow M(v)\cup\{u\}$
30
Algorithm 2 $\varepsilon$-LEDP Matching
4.2 Proof of Theorem 1.3
@@ 4.2.2 Utility
We first prove that the implicitly output $b$-matching is at least half the size of a maximum matching, with high probability.
@@
We will mostly use Lemma 4.2 with $b^{\prime}=1$, but we state and prove the general version needed for Section 10.
@@
For $b=O\left(\frac{\log(n)}{\varepsilon}\right)$, Algorithm 2 outputs an implicit $b$-matching that is at least $\nicefrac{{1}}{{2}}$ the size of a maximum matching, with high probability.
@@
For each iteration $i$, we will show if node $v=v_{i}$ has not satisfied its matching condition and there exists at least $1$ node that comes after it in the order, then it is matched with at least one of the total number of nodes it can be matched with. That is, let $b^{*}(v)\in[0,b]$ denote the maximum number of nodes later than $v$ in the permutation which $v$ can match with (up to $b$); we wish to show that the number of nodes $v$ matches with after iteration $i$ is at least $1$ if $b^{*}(v)\geq 1$. By Lemma 4.2, this implies the utility guarantee. Finally, we show that with high probability, we do not match with more than $b$ nodes.
First, all of the Laplace noise (Algorithms 2, 2, 2 and 2) are drawn from distributions with expectation 00. Let $X$ be a random variable drawn from one of these Laplace distributions. By Lemma A.1, we have that each of the following holds with probability $1-1/n^{3c}$:
@@
$\geq\frac{576c\ln(n)}{\varepsilon}+1-\frac{72c\ln(n)}{\varepsilon^{\prime}}$
$=\frac{576c\ln(n)}{\varepsilon}+1-\frac{72c\ln(n)\cdot(2(1+\eta)+2\cdot\eta)}{\varepsilon\cdot\eta}$ (by our choice of $\varepsilon^{\prime}$ in Algorithm 2)
$\geq\frac{576c\ln(n)}{\varepsilon}+1-\frac{4\cdot 2\cdot 72c\ln(n)}{\varepsilon}$ (by our choice of $\eta=0.5$)
$\geq\frac{576c\ln(n)}{\varepsilon}+1-\frac{576c\ln(n)}{\varepsilon}$
$\geq 1$
where the first line follows since we know that $b\geq\frac{576c\ln(n)}{\varepsilon}+1$. In other words, $v$ was already matched from a previous iteration.
Case II.a: Now, assume that Algorithm 2 does execute. Let $W_{0}(v)$ be the set of unmatched neighbors of $v$ which appear later in the ordering. First consider the case that $|W_{0}(v)|<\frac{8c\ln(n)}{\varepsilon^{\prime}}$. We know that either $|M(v)|=0$ or $|M(v)|\geq 1$. If $|M(v)|\geq 1$, then $v$ is already matched and there is nothing to prove. Otherwise, conditioned on $\mathcal{E}$, $|W_{0}(v)|+3\cdot\frac{12c\ln(n)}{\varepsilon^{\prime}}=\frac{8c\ln(n)}{\varepsilon^{\prime}}+\frac{36c\ln(n)}{\varepsilon^{\prime}}\leq\frac{44c\ln(n)}{\varepsilon^{\prime}}\leq\frac{8\cdot 44c\ln(n)}{\varepsilon}<\frac{576c\ln(n)}{\varepsilon}+1\leq b$ and all of $W_{0}(v)$ is matched. In particular, $v$ is matched to at least $b^{*}(v)$ neighbors.
Case II.b.1: For the remainder of the proof, we assume $k\coloneqq|W_{0}(v)|\geq\frac{8c\ln(n)}{\varepsilon^{\prime}}$. As a reminder, we are in the case that Algorithm 2 does execute. Let $r^{*}$ be the $r$ chosen in Algorithm 2. If the returned $r^{*}>\log_{1+\eta}(k\varepsilon^{\prime}/(8c\ln{n}))$, we know that the condition $|\widetilde{M}(v)|+|\widetilde{W}_{r}(v)|+12c\ln(n)/\varepsilon^{\prime}\leq b$ did not hold for $r=\log_{1+\eta}\left(\frac{k\varepsilon^{\prime}}{8c\ln{n}}\right)$. Thus, we have that $p_{r}=(1+\eta)^{r}=\frac{8c\ln(n)}{k\varepsilon^{\prime}}$ so the expected size of $W_{r}(v)$ is $\frac{8c\ln(n)}{\varepsilon^{\prime}}$; that is, $\mathbb{E}[|W_{r}(v)|]=\frac{8c\ln(n)}{\varepsilon^{\prime}}$. By a multiplicative Chernoff Bound (Theorem A.2), the true size of $W_{r}(v)$ is at most $16c\log(n)/\varepsilon^{\prime}$ with probability at least $1-\frac{1}{n^{4c}}$. Since the condition $|\widetilde{M}(v)|+|\widetilde{W}_{r}(v)|+12c\log(n)/\varepsilon^{\prime}\leq b$ didn’t hold for such $r$, we have that
@@
$|M(v)|$ $>b-\frac{16c\ln(n)}{\varepsilon^{\prime}}-\frac{24c\ln(n)}{\varepsilon^{\prime}}$ ( w.p. at least $1-\frac{1}{n^{4c}}$)
$|M(v)|$ $>b-\frac{40c\ln(n)}{\varepsilon^{\prime}}$
$|M(v)|$ $>b-\frac{4\cdot 40c\ln(n)}{\varepsilon}$ (by the choice of $\varepsilon^{\prime}$)
$|M(v)|$ $>1\,.$ (since $b\geq\frac{576c\ln(n)}{\varepsilon}+1$)
@@
Case II.b.2: If $r^{*}\leq\log_{1+\eta}\left(\frac{k\varepsilon^{\prime}}{8c\ln{n}}\right)$, there are two scenarios. Either $r^{*}=0$, then we match $v$ with all nodes which are still available so clearly the number of nodes $v$ is matched with is at least $b^{*}(v)$, or $r^{*}>1$. In the latter scenario, we know that the threshold is exceeded for $r^{*}-1$, so we have $|\widetilde{M}(v)|+|\widetilde{W}_{r^{*}-1}(v)|+12c\ln(n)/\varepsilon^{\prime}\geq b$. Conditioning on $\mathcal{E}$, $|M(v)|+|{W}_{r^{*}-1}(v)|\geq b-36c\ln(n)/\varepsilon^{\prime}$. Hence, by our assumption that $k\geq\frac{8c\ln(n)}{\varepsilon^{\prime}}$, if $|M(v)|=0$, it holds that $|{W}_{r^{*}-1}(v)|\geq b-36c\ln(n)/\varepsilon^{\prime}$ Thus, by a Chernoff bound, with probability at least $1-\frac{1}{n^{4c}}$
$|W_{r^{*}}(v)|\geq\frac{1-\eta}{1+\eta}\mathbb{E}[|W_{r^{*}-1}(v)|]\geq\frac{1-\eta}{1+\eta}(b-36c\ln(n)/\varepsilon^{\prime})\geq 1\,.$
It follows that $v$ is matched with at least 1 neighbor at the end of this iteration.
Taking a union bound over all $r$ and $v\in V$, all Chernoff bounds hold regardless of the value of $r^{*}$ with probability at least $1-\frac{1}{n^{2c}}$ for $c\geq 3$. Taking into account the conditioning on $\mathcal{E}$, our algorithm succeeds with probability at least $1-\frac{1}{n^{c}}$ for $c\geq 3$.
In summary, if $b^{*}(v)>0$, then $v$ is matched with at least 1 neighbor at the end of the $i$-th iteration. If $b^{*}(v)=0$ and $v$ is not matched, then all of $v$’s neighbors are matched to at least one other node as we showed above. Thus the final output contains a maximal matching as required.
@@
Together, Lemma 4.1 and Lemma 4.3 proves Theorem 1.3.
5 $O(\log n)$-Round $\varepsilon$-LEDP Implicit Matchings

### v3

@@ 4.1 Lower Bound on DP Synthetic Graph Generation
Fix some graph $G\in\mathcal{M}$, and let $uv,ab\in E(G)$. Let $G^{\prime}$ be the graph obtained by removing $uv$ and $ab$, and adding $ub$ and $av$. Note that $G$ and $G^{\prime}$ are distance $4$ away from each other.
For all $y\neq u,b$, note that $uy\notin E(G^{\prime})$. Thus it must be the case that $\mathbb{E}[X_{G^{\prime},uy}^{\mathcal{A}}]\leq b_{u}/(n-2)$, since otherwise the symmetry property of Lemma 4.4 and linearity of expectations implies that the expected (fractional) degree of $u$ in the output of $\mathcal{A}(G)$ is at least $\sum_{y\neq u,b}\mathbb{E}[X_{G^{\prime},uy}^{\mathcal{A}}]>(n-2)\frac{b_{u}}{n-2}=b_{u}$, contradicting our assumption that $\mathcal{A}$ always outputs a feasible solution. Since $G$ and $G^{\prime}$ are at distance $2$ from each other, the fact that $\mathbb{E}[X_{G^{\prime},uv}^{\mathcal{A}}]\leq b_{u}/(n-2)$ implies that
$\mathbb{E}[X_{G,uv}^{\mathcal{A}}]$ $=\int_{0}^{1}\Pr[X_{G,uv}^{\mathcal{A}}\geq x]\mathop{}\!\mathrm{d}x$
$\leq\int_{0}^{1}\left(e^{2\varepsilon}\cdot\Pr[X_{G^{\prime},uv}^{\mathcal{A}}\geq x]+4\delta\right)\mathop{}\!\mathrm{d}x$
$=e^{4\varepsilon}\cdot\mathbb{E}[X_{G^{\prime},uv}^{\mathcal{A}}]+4\delta$
$\leq e^{4\varepsilon}\frac{b_{u}+b_{v}}{2(n-2)}+4\delta\,.$
Here the last inequality uses the fact that if $x\leq y,z$, then $x\leq\frac{1}{2}(y+z)$.
Since this argument applies to all $uv\in E(G)$, we have that
$\mathbb{E}[\nu_{\mathcal{A}}(G)]$ $=\sum_{uv\in E(G)}\mathbb{E}[X_{G,uv}^{\mathcal{A}}]$
$\leq\sum_{uv\in E(G)}\left(e^{4\varepsilon}\frac{b_{u}+b_{v}}{2(n-2)}+4\delta\right)$
$=e^{4\varepsilon}\frac{\bar{b}n}{2(n-2)}+2\delta n.$
$\leq e^{4\varepsilon}\bar{b}+2\delta n.$ ($n\geq 4$)
Since $\nu_{\mathcal{A}}=\min_{G\in\mathcal{M}}\mathbb{E}[\nu_{\mathcal{A}}(G)]$, this implies that $\nu_{\mathcal{A}}\leq e^{4\varepsilon}\bar{b}+2\delta n$, as claimed. ∎
@@
The combination of Lemma 4.4 and Lemma 4.5 imply that any $(\varepsilon,\delta)$-DP algorithm $\mathcal{A}$ which always outputs a fractional stochastic packing solution $X_{G}^{\mathcal{A}}$ satisfies
$\nu_{\mathcal{A}}\coloneqq\sum_{uv\in E(G)}\mathbb{E}[X_{G,uv}^{\mathcal{A}}]\leq e^{4\varepsilon}\bar{b}+2\delta n\,.$
We also know that $|E(G)|=n/2$ for all $G\in\mathcal{M}$. Thus by the definition of an $(\gamma,\beta)$-approximation, it must be the case that $e^{4\varepsilon}\bar{b}+2\delta n\geq\frac{n}{2\gamma}-\beta$. ∎
4.2 Lower Bound on Explicit Matching Solutions (Theorem 2.1)
We first explain how to prove Theorem 2.1 by applying Theorem 4.3. Given $n$ nodes $V$ for an even integer $n>0$, let $\mathcal{M}$ denote the set of all perfect matchings on $V$. Remark that Theorem 4.3 holds even if we restrict the input class to $\mathcal{M}$. As stated in Remark 4.2, over $\mathcal{M}$, any $(\gamma,\beta)$-approximation algorithm for explicit matchings (Definition 3.7) is a $(\gamma,\beta)$-approximation algorithm for the synthetic graph problem (Definition 4.1) as the objective functions coincide and any explicit matching solution is a valid synthetic graph approximation with uniform degree bounds $b_{v}=b$. Thus, Theorem 2.1 follows directly from Theorem 4.3.
5 LEDP Implicit $b$-Matching via Public Vertex Subset Mechanism
@@
In this section, we state our main algorithm for maximal matching (Algorithm 2) and prove its guarantees in Theorem 2.2, which is restated below for readability. This algorithm uses our novel Public Vertex Subset Mechanism (PVSM) which may be applicable to other problems such as vertex cover, dominating set, or independent set. We demonstrate the versatility of PVSM in two different algorithms in this paper. We implement Algorithm 2 for various different models in the later sections. See 2.2
As an immediate corollary, Theorem 2.2 implies the following.
Corollary 5.1 (LEDP $2$-Approximate Maximum Matching).
As mentioned, when $b=\Omega\left(\frac{\log(n)}{\varepsilon}\right)$, Algorithm 2 outputs a $b$-matching that approximates 1-matchings, giving a bicriteria approximation. Moreover, we can avoid a bicriteria approximation when $b$ is slightly larger.
Theorem 5.2 (LEDP $(2+\eta)$-Approximate Maximum $b$-Matching).
For $\varepsilon\in(0,1)$ and $\eta\in(0,1/2)$, Algorithm 2 is an $\varepsilon$-LEDP algorithm that, with high probability, outputs an (implicit) $b$-matching, in the billboard model, for $b=\Omega\left(\frac{\log(n)}{\eta^{3}\varepsilon}\right)$ that has the size of a $(2+\eta)$-approximate maximum $b$-matching.
@@
Our LEDP algorithm draws on a simple procedure for maximal non-private $b$-matching, described as follows. We take an arbitrary ordering on the vertices, and process them one by one. When considering the $i^{th}$ vertex $v_{i}$, let $b^{\prime}$ be the number of additional vertices $v_{i}$ can match with. We wish to find some subset of vertices of size $b^{\prime}$ from later in the ordering with which to match $v_{i}$. Some of these vertices may have already matched with $b$ vertices from previous iterations. We choose an arbitrary subset of $b^{\prime}$ vertices which have not already reached their limit from later in the ordering. We say that vertices which have reached their limit satisfy the matching condition. If there are fewer than $b^{\prime}$ vertices satisfying the matching condition, we match all of them with $v_{i}$. Then we move on to the next vertex in the order. It is clear that this procedure yields a maximal $b$-matching. In the following algorithms, we analyze a suitable privatization of this algorithm and show that given sufficiently large $b$, we are guaranteed to produce an implicit maximal ($1$-)matching. [^7] We remark that all of our algorithms can be straightforwardly adjusted to guarantee maximal $b$-matchings.
To complement our algorithm, we show in Section 5.4 that the bound on $b$ is essentially tight for our specific notion of implicit solution.
Theorem 5.3.
Note that this essentially matches Corollary 5.1 (up to constants) by setting $\beta=1/\mathrm{poly}(n)$.
High-Level Overview.
@@
Organization.
We first give a detailed description of our algorithm, including its pseudocode, in Section 5.1. The accompanying privacy and utility analysis follows in Section 5.2. Section 5.3 proves the special cases of our general result (Theorem 2.2). Finally, we show the complementary lower bound on $b$ in Section 5.4.
@@ 5.1 Detailed Algorithm Description
We now describe our algorithm, Algorithm 2, in detail. The algorithm is provided with a private graph $G=(V,E)$, a privacy parameter $\varepsilon>0$, a matching parameter $b\geq\frac{576c\ln(n)}{\eta\cdot\varepsilon}+1$ and a constant $c\geq 1$ which is used in the high probability guarantee a privacy parameter $\varepsilon>0$, matching benchmark parameter $b^{\prime}\geq 1$, constant $\eta\in(0,1)$, constant $c\geq 3$ which is used in the high probability guarantee, and matching degree parameter $b\geq\frac{(1+\eta)^{2}}{1-\eta}b^{\prime}+\frac{576c\ln(n)}{\eta^{2}\varepsilon}$ . We first set $\varepsilon^{\prime}$ which we use in our algorithm (Algorithm 2). Then, we flip our public set of coins. We flip $O(\log(n))$ coins for each pair of vertices $u\neq v\in V\times V$ (Algorithm 2). Each of the coins flips for pair $\{u,v\}$ is flipped with probability $1/(1+\eta)^{r}$ for each $r\in\{0,\dots,\lceil\log_{1+\eta}(n)\rceil\}$. The result of the coin flip is released and stored in $coin(u,v,r)$ (Algorithm 2).
@@
Input: Graph $G=(V,E)$, privacy parameter $\varepsilon>0$, matching benchmark parameter $b^{\prime}\geq 1$, constant $\eta\in(0,1)$, constant $c\geq 3$, matching degree parameter $b\geq\frac{(1+\eta)^{2}}{1-\eta}b^{\prime}+\frac{576c\ln(n)}{\eta^{2}\varepsilon}$
@@ 14       $u$ releases $A(u)\leftarrow i$
17 if $A(v)>i$ ($v$ has not satisfied its matching condition) then
@@ 27       $M(v)\leftarrow M(v)\cup\{u\}$
Algorithm 2 $\varepsilon$-LEDP Approximate Maximum $b$-Matching
5.2 Analysis (Theorem 2.2)
@@ 5.2.2 Utility
We first prove that the implicitly output $b$-matching is at least half the size of a maximum $b^{\prime}$-matching, with high probability.
@@
The utility proof of Theorem 2.2 is implied by the following lemma, which is a restatement with detailed constants.
@@
First, all of the Laplace noise (Algorithms 2, 2, 2 and 2) are drawn from distributions with expectation 0. Let $X$ be a random variable drawn from one of these Laplace distributions. By Lemma A.1, we have that each of the following holds with probability $1-1/n^{3c}$:
@@
$\geq{b}-\frac{72c\ln(n)}{\varepsilon^{\prime}}$
$\geq{b}-\frac{576c\ln(n)}{\eta\varepsilon}$
$>{b^{\prime}}\,.$ (by $b\geq b^{\prime}+\frac{576c\ln(n)}{\eta\varepsilon}$)
$\leq\frac{238c\ln(n)}{\eta^{2}\varepsilon}+{b^{\prime}}$
$\leq b$
and all of $W_{0}(v)$ is matched.
@@
$|M(v)|$ $>b-\frac{162c\ln(n)}{\eta^{2}\varepsilon}$
@@
with probability $1-\frac{1}{n^{4c}}$.
$\geq b^{\prime}-|M(v)|\,.$
@@
5.3 Corollaries
By choosing a fixed constant $\eta=\nicefrac{{1}}{{2}}$, matching benchmark parameter $b^{\prime}=1$, and matching degree parameter $b=\Omega\left(\frac{\log(n)}{\varepsilon}\right)$, Theorem 2.2 immediately yields a $b$-matching that is the size of a $2$-approximate 11-matching. As stated in Theorem 5.2, when $b=\Omega\left(\frac{\log(n)}{\eta^{3}\varepsilon}\right)$ is slightly larger, we avoid the need for bicriteria approximation and our algorithm yields a $(2+\eta)$-approximate $b$-matching. The proof follows from the following corollary of Lemma 5.6.
Corollary 5.7 ($(2+\eta)$-Approximate Maximum $b$-Matching).
Before proceeding with the proof, we recall that $b$-matchings have the following integral linear programming relaxation [schrijver2003combinatorial, Theorem 33.2], where $\partial(U)$ denotes the set of outgoing edges from $U\subseteq V$.
$\sum_{uv\in E}x(uv)$ $\leq b$ $\forall v\in V$
$x$ $\in[0,1]^{E}$
Proof.
Choose $b^{\prime}$ such that
5.4 Lower Bound for Implicit Matchings
We now prove Theorem 5.3, our lower bound for implicit solutions, which essentially matches our upper bound from Corollary 5.1. While it does not apply to all algorithms in the billboard model, it applies to the implicit solutions that our algorithms use (Definition 3.6).
Recall that, given an input graph $H$, our implicit solutions essentially output a subset of vertices for each node $x$ whose intersection with $x$’s private adjacency list gives the nodes $x$ is matched to. We denote these public subsets of vertices by $\mathcal{S}=\{S_{x}\}_{x\in V}$ where $S_{x}$ is the public subset of vertices for node $x$. That is, $\mathcal{S}=\{S_{x}\}_{x\in V}$ is the implicit solution generated by $\mathcal{A}$. We say that $\mathcal{A}$ includes an edge $\{u,v\}$ if either $u\in S_{v}$ or $v\in S_{u}$. In other words, $\mathcal{A}$ includes an edge if that edge is an edge in $H(\mathcal{S})$.
Proof of Theorem 5.3.
[…]
6 $O(\log n)$-Round LEDP Implicit $b$-Matching via PVSM